# Quantile Reward Policy Optimization: Alignment with Pointwise Regression and Exact Partition Functions

### v1

@@ Appendix I Noise amplitude in the transformed reward estimate
SFT $\rightarrow$ offpolicy–best* ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}\underline{0.1469}}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0003}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1480}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0007}}$ ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}\underline{0.1459}}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0003}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1465}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0008}}$ 0.03 1e-6 100
SFT $\rightarrow$ offpolicy–random ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1441}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0001}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1454}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0004}}$ ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1430}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0003}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1440}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0006}}$ 0.01 3e-7 476
@@
offpolicy–best ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1433}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0003}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1450}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0003}}$ ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1443}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0005}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1450}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0023}}$ 10 3e-7 200
offpolicy–random ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1419}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0003}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1431}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0004}}$ ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1419}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0005}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1432}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0017}}$ 2.5 3e-7 476
@@
SFT $\rightarrow$ offpolicy–best* ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}\underline{0.1460}}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0002}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1472}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0005}}$ ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}\underline{0.1467}}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0003}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1478}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0007}}$ 10 3e-7 200
SFT $\rightarrow$ offpolicy–random ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1435}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0005}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1441}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0003}}$ ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1445}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0005}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1447}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0004}}$ 10 3e-7 476

### v2

@@ Appendix I Noise amplitude in the partition function estimate for the initial reward
SFT $\rightarrow$ off-policy–best* ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}\underline{0.1469}}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0003}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1480}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0007}}$ ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}\underline{0.1459}}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0003}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1465}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0008}}$ 0.03 1e-6 100
SFT $\rightarrow$ off-policy–random ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1441}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0001}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1454}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0004}}$ ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1430}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0003}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1440}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0006}}$ 0.01 3e-7 476
@@
off-policy–best ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1433}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0003}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1450}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0003}}$ ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1443}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0005}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1450}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0023}}$ 10 3e-7 200
off-policy–random ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1419}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0003}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1431}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0004}}$ ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1419}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0005}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1432}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0017}}$ 2.5 3e-7 476
@@
SFT $\rightarrow$ off-policy–best* ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}\underline{0.1460}}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0002}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1472}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0005}}$ ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}\underline{0.1467}}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0003}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1478}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0007}}$ 10 3e-7 200
SFT $\rightarrow$ off-policy–random ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1435}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0005}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1441}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0003}}$ ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1445}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0005}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1447}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0004}}$ 10 3e-7 476